### v1

@@ Abstract
In this paper, the light-front wave functions (LFWFs) of $\rho$ meson are evaluated in the framework of Nambu–Jona-Lasinio model using the proper time regularization scheme. The transverse momentum dependent parton distributions (TMDs) of the $\rho$ meson are derived from the overlap representations of the LFWFs. We investigate the $\bm{k}_{\perp}$-weighted moments and the $x$-dependent average transverse momentum $\langle k_{\perp}^{n}(x)\rangle_{\alpha}$ of $\rho$ meson TMDs, $\langle k_{\perp}^{n}(x)\rangle_{\alpha}$ shows the typical transverse momenta of quark TMDs in our model. Our findings indicate that the average transverse momenta of $\langle k_{\perp}(x)\rangle_{\alpha}$ fall within the range of $[0.3,0.45]$ GeV, while $\langle k_{\perp}^{2}(x)\rangle_{\alpha}$ are in the region of $[0.15,0.30]$ GeV. Our TMDs and parton distribution function (PDFs) exhibit excellent adherence to positive constraints. The investigation of spin densities within the $\rho$ meson in transverse momentum space reveals axially symmetric distributions for quarks and targets polarized in either longitudinal or transverse directions. Conversely, unpolarized symmetric distributions occur when the quark is longitudinally polarized and the target is transversely polarized, while exhibiting dipolar distortions in the opposite scenario. A comparison between TMDs from LFWFs and those from a covariant approach demonstrates that TMDs from LFWFs more effectively satisfy positive constraints than their counterparts from the covariant approach. Furthermore, the $x$-moments of $\rho$ meson PDFs indicate that $\langle x\rangle_{g_{L}}$ is larger using LFWFs compared to the covariant approach. Additionally, analysis of quark spin densities within the $\rho$ meson reveals that $\rho_{\uparrow\downarrow}\left(k_{x},k_{y}\right)=f-g_{L}-\frac{1}{3}f_{LL}$ yields a negative result using the covariant approach, whereas $\rho_{\uparrow\downarrow}\left(k_{x},k_{y}\right)$ derived from LFWFs produces a positive outcome.
@@ I Introduction
The most difficult job of hadron physics is to compute the hadronic transition matrix elements which are fully governed by non-perturbative Quantum Chromodynamics (QCD). Due to the lack of complete understanding of the non-perturbative QCD so far, one needs to utilize phenomenological models. In describing various processes, such models have achieved relative successes up to now. Among these models the Nambu–Jona-Lasinio (NJL) model Klevansky (1992); Buballa (2005); Cloët et al. (2014); Cui et al. (2017, 2016); Zhang et al. (2016, 2018); Zhang and Wu (2024) is a Poincaré covariant quantum field theory with many of the same low-energy properties as QCD. It has the same flavor symmetries as QCD and should therefore offer a stable chiral effective theory of QCD valid at low to intermediate energies. Above all, it includes the key prominent phenomena of confinement and dynamical chiral symmetry breaking.
@@
A variety of high energy hadronic processes and experimentally observable quantities can be interpreted in terms of LFWFs, for example, electromagnetic form factors (FFs) Bartel et al. (1973); Hohler et al. (1976); Vogel and Engel (1989); Mergell et al. (1996); Kelly (2004); Perdrisat et al. (2007); Puckett et al. (2010); De Mori (2022); Stamen et al. (2022); Batelaan et al. (2022) and PDFs Holt and Roberts (2010); Nguyen et al. (2011); Dulat et al. (2016); Ding et al. (2020); Cui et al. (2021, 2020) can be analyzed using LFWFs. In addition, LFWFs also can be used to estimate of weak decay rates Brodsky and Hwang (1999); Choi and Ji (2007), calculate single spin asymmetries in semi-inclusive deep inelastic scattering Brodsky et al. (2002); Gamberg et al. (2008), and determine the cross sections of electromagnetic production for vector mesons Radyushkin (1996); Vanderhaeghen et al. (1999); Goloskokov and Kroll (2007). Furthermore, the three-dimensional structure of the nucleon in the momentum and coordinate space characterized by the TMDs Collins et al. (2008); Meissner et al. (2007); Hautmann (2007); Barone et al. (2010); Puhan and Dahiya (2024); Xie and Lu (2023); Puhan et al. (2024) and generalized parton distributions (GPDs) Müller et al. (1994); Ji (1997); Radyushkin (1997); Vanderhaeghen et al. (1999); Belitsky et al. (2002); Goeke et al. (2001); Burkardt (2003); Diehl (2003); Belitsky and Radyushkin (2005); Zhang et al. (2021b, c, 2022a); Adhikari et al. (2021); Fu et al. (2022); Zhang et al. (2022b); Luan and Lu (2024). TMDs are sensitive to spin dependent correlations and the parton’s intrinsic transverse motion, offering a more complete understanding of the parton structures, especially the transverse structures of hadrons.
@@
In this paper, we have computed the leading Fock-state LFWFs of the $\rho$ meson within the framework of the Nambu-Jona-Lasinio (NJL) model. Specifically, we focus on the $\rho$ meson’s $|q\bar{q}\rangle$ component. The NJL model is characterized by an effective Lagrangian for relativistic fermions interacting through local fermion-fermion couplings, and it finds widespread application in various fields.
@@ II.1 NJL Model
The SU(2) flavor NJL Lagrangian is
@@
where $\hat{m}=\text{diag}\left(m_{u},m_{d}\right)$ is current quark mass matrix and $\vec{\tau}$ are the Pauli matrices to stand for isospin. In the limit of exact isospin symmetry, $m_{u}=m_{d}=m$. The 4-fermion coupling constants in chiral channels are labeled by $G_{\pi}$ , $G_{\omega}$, and $G_{\rho}$.
The dressed quark propagator is,
@@
The interaction kernel of the gap equation is local, one can obtain a constant dressed quark mass $M$, which satisfies
$M=m+12iG_{\pi}\int\frac{\mathrm{d}^{4}l}{(2\pi)^{4}}\text{tr}_{\text{D}}[S(l)],$ (3)
where the trace is over Dirac indices.
NJL model is not renormalizable, it needs a regularization scheme to fully define the model. We will use the proper time regularization scheme Ebert et al. (1996); Hellstern et al. (1997); Bentz and Thomas (2001).
@@
where $X$ stands for a product of propagators that have been combined using Feynman parametrization. We introduce the infrared cutoff $\Lambda_{\text{IR}}$ to mimic confinement because it is absent in the NJL model. The infrared cutoff should be of the order $\Lambda_{\text{QCD}}$ and we choose $\Lambda_{\text{IR}}=0.240$ GeV. The coupling strength $G_{\pi}$, the momentum cutoff $\Lambda_{\text{UV}}$ and the current quark mass $m$ are determined through the Gell-Mann–Oakes–Renner (GMOR) relation, $f_{\pi}^{2}m_{\pi}^{2}=-m\langle\bar{\psi}\psi\rangle$ and gap equation $M=m-2G_{\pi}\langle\bar{\psi}\psi\rangle$, where $\langle\bar{\psi}\psi\rangle$ is two-quark condensate derived from QCD sum rules. $m_{\pi}=0.140$ GeV is the physical pion mass, $f_{\pi}=0.092$ GeV is the pion decay constant, $m$ and $M$ are the current and the constituent quark mass, respectively.
@@
where $\bar{x}=1-x$. $m_{\pi}=0.14$ GeV is coinciding with the value obtained from the pole condition $01+2G_{\pi}\Pi_{PP}(-m_{\pi}^{2})=$ 0. $G_{\omega}$ and $G_{\rho}$ are determined by $m_{\omega}=0.782$ GeV, $m_{\rho}=0.770$ GeV through $01+2G_{i}\Pi_{VV}(-m_{i}^{2})=$ 0, $i=(\omega,\rho)$, where $\Pi_{VV}(Q^{2})$ is the self-energy of the vector bubble diagram defined in Eq. (7). The parameters used in this work are listed in Table 1.
From the inhomogeneous Bethe-Salpeter equation, the dressed quark form factors, associated with the electromagnetic currents, are Zhang et al. (2021a)
@@ II.2 The definition and calculation of LFWFs
where $a^{0}=\frac{1}{\sqrt{2}}(a^{0}+a^{3})$, $\bm{a}_{\perp}=(a_{1},a_{2})$ are the light-cone $\pm$ and transverse components of a 4-vector $a^{\mu}$. The Lorentz invariant spin structure for the vector meson is expressed by accounting the photon-quark-antiquark vertex:
@@
then we obtain the final form of
@@ III.1 Overlap formalism
Figure 1: Results for the quark TMDs of $\rho$ meson as a function of $x$ and $\bm{k}_{\perp}^{2}$. The TMDs are given in units of GeV-2 and $\bm{k}_{\perp}^{2}$ in units of GeV2.
The explicit overlap form for all the T-even TMDs in terms of helicity amplitudes is expressed as Shi et al. (2022)
@@ III.2 Transverse momentum dependence
$\langle k_{\perp}^{2}(x)\rangle_{f}$ $\approx 0.25\quad\text{GeV},$ (30)
@@
$\langle k_{\perp}^{2}(x)\rangle_{g_{T}}$ $=\langle k_{\perp}^{2}(x)\rangle_{h}\simeq[0.16,0.24]\quad\text{GeV},$ (32)
@@
$\langle k_{\perp}^{2}(x)\rangle_{f_{TT}}$ $=\langle k_{\perp}^{2}(x)\rangle_{h_{L}^{\perp}}\simeq[0.14,0.19]\quad\text{GeV}.$ (34)
@@
$\langle k_{\perp}^{2}(x)\rangle_{g_{L}}$ $\in[0.10,0.24]\quad\text{GeV},$ (39)
@@ III.3 PDFs
$=\frac{3N_{L}^{2}}{4\pi^{2}}\int\mathrm{d}\tau\frac{1}{\tau}(M^{2}+\frac{1}{\tau})e^{-\tau B}$
@@
Figure 4: The upper panel: the unpolarized, tensor polarized, helicity and transversity PDFs of $\rho$ meson in this paper. The lower panel: the unpolarized, tensor polarized, helicity and transversity PDFs of $\rho$ meson in Ref. Ninomiya et al. (2017).
@@
which means the total contribution of the valence quark and antiquark to the spin of the $\rho$ meson is $50.5\%$, indicating that approximately $49.5\%$ of the contribution comes from the quark orbital angular momentum. In contrast, Ref. Ninomiya et al. (2017) reports this value as $56\%$, which is higher.
@@ IV Summary and conclusion
We study the transverse momentum dependence of the $\rho$ meson TMDs, with a focus on their $\bm{k}_{\perp}$-weighted moments. In Ref. Ninomiya et al. (2017), the authors also calculate the $\rho$ meson TMDs in the NJL model using the covariant method. A comparison of $k_{\perp}$-weighted moments with their results shows that the numerical values are almost identical, except for $\langle k_{\perp}^{n}\rangle_{g_{L}}$, where our values are larger than theirs. Notably, our numerical values of $\langle k_{\perp}^{n}\rangle_{g_{L}}$ align with those of $\langle k_{\perp}^{n}\rangle_{h_{L}^{\perp}}$, consistent with findings in Refs. Kaur et al. (2021); Shi et al. (2022). Furthermore, we investigate the $x$-dependent average transverse momentum $\langle k_{\perp}^{n}(x)\rangle_{\alpha}$ of TMDs to determine typical transverse momenta of quark TMDs. Our results indicate that the typical transverse momenta of $\langle k_{\perp}(x)\rangle_{\alpha}$ fall within the range $[0.3,0.45]$ GeV and those of $\langle k_{\perp}^{2}(x)\rangle_{\alpha}$ lie within $[0.15,0.30]$ GeV region.
The four $\rho$ meson PDFs can be obtained by integrating the TMDs over $\bm{k}_{\perp}$. A comparison with Ref. Ninomiya et al. (2017) reveals that $f(x)$, $f_{LL}(x)$, $h(x)$ and $h(x)$ are similar, but there is a difference in $g(x)$. In Ref. Ninomiya et al. (2017), $g(x)$ is negative when $x$ is small, whereas our results show that it is positive in the entire region of $x\in[0,1]$. We also conducted an investigation into the $x$-moments of various PDFs. Our spin sum rule indicates that $\Delta q=\langle x\rangle_{g}=0.505$, suggesting that the total valence quark and antiquark contribution to the spin of the $\rho$ meson is $50.5\%$, implying that about $49.5\%$ contribution comes from quark orbital angular momentum. The $x$-moments of $h(x)$ yield $\langle x\rangle_{h}=0.803$, indicating that the tensor charge at the model scale is 0.803.

### v3

@@ Abstract
In this paper, the light-front wave functions (LFWFs) of $\rho$ meson are evaluated in the framework of Nambu–Jona-Lasinio model using the proper time regularization scheme. The transverse momentum dependent parton distributions (TMDs) of the $\rho$ meson are derived from the overlap representations of the LFWFs. We investigate the $\bm{k}_{\perp}$-weighted moments and the $x$-dependent average transverse momentum $\langle k_{\perp}^{n}(x)\rangle_{\alpha}$ of $\rho$ meson TMDs, $\langle k_{\perp}^{n}(x)\rangle_{\alpha}$ shows the typical transverse momenta of quark TMDs in our model. Our findings indicate that the average transverse momenta of $\langle k_{\perp}(x)\rangle_{\alpha}$ fall within the range of $[0.3,0.45]$ GeV, while $\langle k_{\perp}^{2}(x)\rangle_{\alpha}$ are in the region of $[0.15,0.30]$ GeV2. Our TMDs and parton distribution function (PDFs) exhibit excellent adherence to positive constraints. The investigation of spin densities within the $\rho$ meson in transverse momentum space reveals axially symmetric distributions for quarks and targets polarized in either longitudinal or transverse directions. Conversely, unpolarized symmetric distributions occur when the quark is longitudinally polarized and the target is transversely polarized, while exhibiting dipolar distortions in the opposite scenario. A comparison between TMDs from LFWFs and those from a covariant approach demonstrates that TMDs from LFWFs more effectively satisfy positive constraints than their counterparts from the covariant approach. Furthermore, the $x$-moments of $\rho$ meson PDFs indicate that $\langle x\rangle_{g_{L}}$ is larger using LFWFs compared to the covariant approach. Additionally, analysis of quark spin densities within the $\rho$ meson reveals that $\rho_{\uparrow\downarrow}\left(k_{x},k_{y}\right)=f-g_{L}-\frac{1}{3}f_{LL}$ yields a negative result using the covariant approach, whereas $\rho_{\uparrow\downarrow}\left(k_{x},k_{y}\right)$ derived from LFWFs produces a positive outcome.
@@ I Introduction
The most difficult job of hadron physics is to compute the hadronic transition matrix elements which are fully governed by non-perturbative Quantum Chromodynamics (QCD). Due to the lack of complete understanding of the non-perturbative QCD so far, one needs to utilize phenomenological models. In describing various processes, such models have achieved relative successes up to now. Among these models the Nambu–Jona-Lasinio (NJL) model Klevansky (1992); Buballa (2005); Cloët et al. (2014); Cui et al. (2017, 2016); Zhang et al. (2016, 2018); Zhang and Wu (2024); Zhang (2024) is a Poincaré covariant quantum field theory with many of the same low-energy properties as QCD. It has the same flavor symmetries as QCD and should therefore offer a stable chiral effective theory of QCD valid at low to intermediate energies. Above all, it includes the key prominent phenomena of confinement and dynamical chiral symmetry breaking.
@@
A variety of high energy hadronic processes and experimentally observable quantities can be interpreted in terms of LFWFs, for example, electromagnetic form factors Bartel et al. (1973); Hohler et al. (1976); Vogel and Engel (1989); Mergell et al. (1996); Kelly (2004); Perdrisat et al. (2007); Puckett et al. (2010); De Mori (2022); Stamen et al. (2022); Batelaan et al. (2022) and PDFs Holt and Roberts (2010); Nguyen et al. (2011); Dulat et al. (2016); Ding et al. (2020); Cui et al. (2021, 2020) can be analyzed using LFWFs. In addition, LFWFs also can be used to estimate of weak decay rates Brodsky and Hwang (1999); Choi and Ji (2007), calculate single spin asymmetries in semi-inclusive deep inelastic scattering Brodsky et al. (2002); Gamberg et al. (2008), and determine the cross sections of electromagnetic production for vector mesons Radyushkin (1996); Vanderhaeghen et al. (1999); Goloskokov and Kroll (2007). Furthermore, the three-dimensional structure of the nucleon in the momentum and coordinate space characterized by the TMDs Collins et al. (2008); Meissner et al. (2007); Hautmann (2007); Barone et al. (2010); Puhan and Dahiya (2024); Xie and Lu (2023); Puhan et al. (2024) and generalized parton distributions (GPDs) Müller et al. (1994); Ji (1997); Radyushkin (1997); Vanderhaeghen et al. (1999); Belitsky et al. (2002); Goeke et al. (2001); Burkardt (2003); Diehl (2003); Belitsky and Radyushkin (2005); Zhang et al. (2021b, c, 2022a); Adhikari et al. (2021); Fu et al. (2022); Zhang et al. (2022b); Luan and Lu (2024). TMDs are sensitive to spin dependent correlations and the parton’s intrinsic transverse motion, offering a more complete understanding of the parton structures, especially the transverse structures of hadrons.
@@
In this paper, we have computed the leading Fock-state LFWFs of the $\rho$ meson within the framework of the NJL model. The NJL model is characterized by an effective Lagrangian for relativistic fermions interacting through local fermion-fermion couplings, and it finds widespread application in various fields. In the $\rho$ meson channel, the model is equivalent to the symmetry-preserving treatment of a contact interaction Gutierrez-Guerrero et al. (2010); Roberts et al. (2011); hence, may be seen as an application of continuum Schwinger function methods Ding et al. (2023). Viewed from this perspective, our calculation may be understood as an analysis of $\rho$ meson properties at the hadron scale, $\zeta_{\cal H}$, called $Q_{0}$ herein, whereat dressed quark and antiquark degrees-of-freedom carry all hadron properties. We do not consider evolution to resolving scales $Q>Q_{0}$; but that may be accomplished using the all-orders evolution scheme detailed elsewhere Yin et al. (2023).
@@ II.1 NJL Model
The NJL model focuses on quarks as the explicit degrees of freedom, with gluons only expressed implicitly via the infrared strength of the contact interaction. The Lagrangian for SU(2) flavor in the $qq$ channel is given by
@@
where $\psi$ is the quark field, $\hat{m}\equiv\text{diag}\left(m_{u},m_{d}\right)$ is current quark mass matrix and in the isospin symmetry $m_{u}=m_{d}=m$. $\vec{\tau}$ are the Pauli matrices represent isospin. In chiral channels the 4-fermion coupling constants are labeled by $G_{\pi}$ , $G_{\omega}$, and $G_{\rho}$. In working with Eq. (II.1), our calculations deliver properties of what will emerge as the dynamically generated dressed-quark + dressed-antiquark core of the $\rho$ meson. This is known to be a good approximation – see, e.g., Ref. Roberts et al. (2011).
In the NJL model, the solution for the dressed quark propagator is,
@@
A regularization scheme is needed to fully define the model because NJL model is not renormalizable. We will implement the proper time regularization (PTR) scheme Ebert et al. (1996); Hellstern et al. (1997); Bentz and Thomas (2001).
@@
where $X$ represents a product of propagators that have been combined using Feynman parametrization. The infrared cutoff should be approximately at the scale of $\Lambda_{\text{QCD}}$, and we have selected $\Lambda_{\text{IR}}=0.240$ GeV, as it incorporates the confinement scale into our calculations. The coupling strength $G_{\pi}$, the momentum cutoff $\Lambda_{\text{UV}}$, and the current quark mass $m$ are determined through the Gell-Mann–Oakes–Renner (GMOR) relation, which is expressed as $f_{\pi}^{2}m_{\pi}^{2}=-m\langle\bar{\psi}\psi\rangle$. Additionally, the gap equation $M=m-2G_{\pi}\langle\bar{\psi}\psi\rangle$ is utilized in this determination. $\langle\bar{\psi}\psi\rangle$ represents a two-quark condensate derived from QCD sum rules. Here, $m_{\pi}=0.140$ GeV represents the physical pion mass, while $f_{\pi}=0.092$ GeV denotes the pion decay constant. $m$ and $M$ refer to the current and constituent quark masses, respectively.
@@
where $\bar{x}=1-x$, and $m_{\pi}=0.14$ GeV coincides with the value obtained from the pole condition $01+2G_{\pi}\Pi_{PP}(-m_{\pi}^{2})=$ 0. The values of $G_{\omega}$ and $G_{\rho}$ are determined by $m_{\omega}=0.782$ GeV, $m_{\rho}=0.770$ GeV through the equation $01+2G_{i}\Pi_{VV}(-m_{i}^{2})=$ 0, where $i=(\omega,\rho)$. Here, $\Pi_{VV}(Q^{2})$ represents the self-energy of the vector bubble diagram as defined in Eq. (6). The parameters utilized in this study are listed in Table 1.
The absence of an explicit $Q^{2}$ scale is a common limitation in most model determinations of quark distributions. As a result, the model scale, denoted as $Q_{0}^{2}$, must be established through comparison with empirical data. We adopt $Q_{0}^{2}=0.16$ GeV2, consistent with Ref. Cloet et al. (2008), which is representative of models dominated by valence contributions Schreiber et al. (1991); Mineo et al. (1999, 2002); Ninomiya et al. (2017).
The form factors of the dressed quark associated with electromagnetic currents are derived from the inhomogeneous Bethe-Salpeter equation Zhang et al. (2021a)
@@ II.2 The definition and calculation of LFWFs
where $a^{\pm}=\frac{1}{\sqrt{2}}(a^{0}\pm a^{3})$, $\bm{a}_{\perp}=(a_{1},a_{2})$ are the light-cone $\pm$ and transverse components of a 4-vector $a^{\mu}$. The Lorentz invariant spin structure for the vector meson is expressed by accounting the photon-quark-antiquark vertex:
@@
then we obtain the final form of the LFWFs
@@ III.1 Overlap formalism
Figure 1: Results for the quark TMDs of $\rho$ meson as a function of $x$ and $\bm{k}_{\perp}^{2}$ at the model scale $Q_{0}^{2}=0.16$ GeV2. The TMDs are given in units of GeV-2 and $\bm{k}_{\perp}^{2}$ in units of GeV2.
The explicit overlap form for all the T-even TMDs in terms of helicity amplitudes are expressed as Shi et al. (2022)
@@ III.2 Transverse momentum dependence
$\langle k_{\perp}^{2}(x)\rangle_{f}$ $\approx 0.25\quad\text{GeV}^{2},$ (29)
@@
$\langle k_{\perp}^{2}(x)\rangle_{g_{T}}$ $=\langle k_{\perp}^{2}(x)\rangle_{h}\simeq[0.16,0.24]\quad\text{GeV}^{2},$ (31)
@@
$\langle k_{\perp}^{2}(x)\rangle_{f_{TT}}$ $=\langle k_{\perp}^{2}(x)\rangle_{h_{L}^{\perp}}\simeq[0.14,0.19]\quad\text{GeV}^{2}.$ (33)
@@
$\langle k_{\perp}^{2}(x)\rangle_{g_{L}}$ $\in[0.10,0.24]\quad\text{GeV}^{2},$ (38)
@@ III.3 PDFs
$+\frac{3N_{L}^{2}}{4\pi^{2}}\int\mathrm{d}\tau\frac{1}{\tau}(M^{2}+\frac{1}{\tau})e^{-\tau B}$
@@
Figure 4: The unpolarized, tensor-polarized, helicity, and transversity PDFs of the $\rho$ meson presented in this paper at the model scale $Q_{0}^{2}=0.16$ GeV2 are compared with the results obtained by Yu Ninomiya in Ref. Ninomiya et al. (2017) at the same model scale.
@@
which means the total contribution of the valence quark and antiquark to the spin of the $\rho$ meson is $50.5\%$. In contrast, Ref. Ninomiya et al. (2017) reports this value as $56\%$, which is higher.
@@ IV Summary and conclusion
We study the transverse momentum dependence of the $\rho$ meson TMDs, with a focus on their $\bm{k}_{\perp}$-weighted moments. In Ref. Ninomiya et al. (2017), the authors also calculate the $\rho$ meson TMDs in the NJL model using the covariant method. A comparison of $k_{\perp}$-weighted moments with their results shows that the numerical values are almost identical, except for $\langle k_{\perp}^{n}\rangle_{g_{L}}$, where our values are larger than theirs. Notably, our numerical values of $\langle k_{\perp}^{n}\rangle_{g_{L}}$ align with those of $\langle k_{\perp}^{n}\rangle_{h_{L}^{\perp}}$, consistent with findings in Refs. Kaur et al. (2021); Shi et al. (2022). Furthermore, we investigate the $x$-dependent average transverse momentum $\langle k_{\perp}^{n}(x)\rangle_{\alpha}$ of TMDs to determine typical transverse momenta of quark TMDs. Our results indicate that the typical transverse momenta of $\langle k_{\perp}(x)\rangle_{\alpha}$ fall within the range $[0.3,0.45]$ GeV and those of $\langle k_{\perp}^{2}(x)\rangle_{\alpha}$ lie within $[0.15,0.30]$ GeV2 region.
The four $\rho$ meson parton distribution function (PDFs) can be obtained by integrating the TMDs over $\bm{k}_{\perp}$. A comparison with Ref. Ninomiya et al. (2017) reveals that $f(x)$, $f_{LL}(x)$, $h(x)$ and $h(x)$ are similar, but there is a difference in $g(x)$. In Ref. Ninomiya et al. (2017), $g(x)$ is negative when $x$ is small, whereas our results show that it is positive in the entire region of $x\in[0,1]$. We also conducted an investigation into the $x$-moments of various PDFs. Our spin sum rule indicates that $\Delta q=\langle x\rangle_{g}=0.505$, suggesting that the total valence quark and antiquark contribution to the spin of the $\rho$ meson is $50.5\%$, implying that about $49.5\%$ contribution comes from quark orbital angular momentum. The $x$-moments of $h(x)$ yield $\langle x\rangle_{h}=0.803$, indicating that the tensor charge at the model scale is 0.803.